### v1

# miniF2F-Dafny: LLM-Guided Mathematical Theorem Proving via Auto-Active Verification

Abstract.
We present miniF2F-Dafny, the first translation of the mathematical reasoning benchmark miniF2F to an automated theorem prover: Dafny. Previously, the benchmark existed only in interactive theorem provers (Lean, Isabelle, HOL Light, Metamath). We find that Dafny’s automation verifies 99/244 (40.6%) of the test set and 109/244 (44.7%) of the validation set with empty proofs—requiring no manual proof steps. For problems where empty proofs fail, we evaluate 12 off-the-shelf LLMs on providing proof hints. The best model we test achieves 55.7% pass@4 success rate employing iterative error correction. These preliminary results highlight an effective division of labor: LLMs provide high-level guidance while automation handles low-level details. Our benchmark can be found on GitHub.
copyright: noneconference: ; 11th January, 2026; Rennes, France
include "../definitions.dfy"
include "../library.dfy"
lemma imo_1959_p1(n: nat)
requires 0 < n
ensures gcd(21*n + 4, 14*n + 3) == 1
(a) Dafny (no LLM)
import Mathlib
theorem imo_1959_p1
(n : ℕ)
(h : 0 < n) :
Nat.gcd (21*n + 4) (14*n + 3) = 1 := by
have eq1 : (21 : ℕ) * n + (4 : ℕ) = ((7 : ℕ) * n + (1 : ℕ)) + ((14 : ℕ) * n + (3 : ℕ)) * (1 : ℕ) := by ring
conv =>
lhs
arg 1
rw [eq1]
rw [Nat.gcd_add_mul_left_left ((7 : ℕ) * n + (1 : ℕ)) ((14 : ℕ) * n + (3 : ℕ)) (1 : ℕ)]
rw [Nat.gcd_comm]
have eq2 : (14 : ℕ) * n + (3 : ℕ) = (1 : ℕ) + ((7 : ℕ) * n + (1 : ℕ)) * (2 : ℕ) := by ring
conv =>
lhs
arg 1
rw [eq2]
rw [Nat.gcd_add_mul_left_left (1 : ℕ) ((7 : ℕ) * n + (1 : ℕ)) (2 : ℕ)]
rw [Nat.gcd_one_left ((7 : ℕ) * n + (1 : ℕ))]
(b) Lean (SeedProver(seed-prover))
Figure 1. Dafny and Lean proofs for IMO 1959 Problem 1.
1. Introduction
Formal mathematical reasoning has long been considered a hallmark of human intelligence, requiring both creative insight and rigorous logical deduction. The advent of large language models (LLMs) has opened up new potential avenues in automating mathematical formalization—the process of translating informal mathematical statements and proofs into machine-verifiable formal systems. While early work focused on interactive theorem provers (ITPs) like Lean, Isabelle, and Coq, these systems require extensive manual proof construction, where human experts or AI systems must provide explicit proof terms for every logical step.
To evaluate progress in AI-assisted formal theorem proving, several benchmarks have been developed. The miniF2F benchmark, originally developed for Lean, consists of 488 mathematical problems drawn from the AIME, AMC, and IMO, as well as material from high school and undergraduate mathematics courses, split into a test set of 244 problems and a validation set of 244 problems. It has become a standard evaluation benchmark for AI systems attempting formal mathematical reasoning. However, all existing translations (Lean, Isabelle, HOL Light, Metamath) target interactive theorem provers, where automation is limited and proofs require significant manual effort. While recent approaches have achieved strong results on miniF2F in Lean, they rely on either specialized models trained extensively on formal proof corpora or elaborate agentic frameworks that orchestrate multiple components to generate the detailed proof scripts that interactive theorem provers require.
We introduce miniF2F-Dafny, the first translation of the miniF2F benchmark to an auto-active verification language. Unlike interactive theorem provers, Dafny is built around automated reasoning capabilities powered by SMT solvers, offering a fundamentally different paradigm: many mathematical problems can be verified with minimal or no manual proof steps. This raises the question of whether the reduced proof complexity in auto-active verifiers enables off-the-shelf LLMs to effectively provide high-level mathematical insights—such as key lemmas, proof strategies, or intermediate assertions—while delegating low-level logical reasoning to automated solvers.
We find that Dafny’s automation alone solves 40-45% of miniF2F problems with empty proofs, requiring zero human input. For the remaining problems, we evaluate 12 off-the-shelf LLMs on providing proof hints to guide Dafny’s automation. Our results highlight an effective division of labor: instead of generating complete proofs, LLMs provide high-level guidance while automation handles low-level details.
Our main contributions are as follows:
  • We present miniF2F-Dafny, the first translation of miniF2F to an auto-active verification language.
  • We establish baseline results showing that Dafny’s automation alone solves 40-45% of problems with no manual intervention.
  • We evaluate 12 off-the-shelf LLMs on providing proof hints for problems requiring human guidance, achieving 55.7% [^1] In Table 1, # denotes problems attempted, including 99/244 problems solved by the Dafny verifier baseline. success rate with our best model.
  • We demonstrate the value of a division of labor in AI-assisted theorem proving: automated solvers handle low-level logical steps while LLMs focus on high-level proof structure.
Model # [^1] In Table 1, # denotes problems attempted, including 99/244 problems solved by the Dafny verifier baseline. Pass@1222In Table 1, Pass@N indicates projected pass rate with up to N generation attempts for models with partial evaluation progress. @2222In Table 1, Pass@N indicates projected pass rate with up to N generation attempts for models with partial evaluation progress. @4222In Table 1, Pass@N indicates projected pass rate with up to N generation attempts for models with partial evaluation progress.
Claude Sonnet 4.5 244 52.5 53.7 55.7
Claude Sonnet 4 243 49.7 50.9 53.0
Claude Sonnet 3.7 164 48.8 52.5 55.2
Claude Haiku 4.5 244 42.2 43.9 43.9
DeepSeek V3.1 222 44.0 45.9 47.3
Llama 4 Maverick 17B 183 44.1 44.1 44.1
GPT-OSS 120B 191 42.5 43.8 45.1
GPT-OSS 20B 244 41.4 42.2 43.4
Qwen 3 235B A22B 164 47.0 53.4 54.3
Qwen 3 Coder 480B A35B 244 45.5 47.5 48.8
Qwen 3 32B 244 43.0 43.9 45.5
Qwen 3 Coder 30B 244 43.0 44.3 44.7
Dafny Verifier (No LLM) 244 40.6 40.6 40.6
Table 1. Partial evaluation results on the miniF2F-Dafny test set.
2. miniF2F-Dafny Benchmark
2.1. Overview
The miniF2F-Dafny benchmark consists of 488 mathematical problems (244 test, 244 validation) translated from the Lean miniF2F benchmark. Each problem is formulated as a Dafny lemma with preconditions (requires clauses) and postconditions (ensures clauses), but an empty proof body. The task is to synthesize a proof that satisfies Dafny’s verifier when the empty proof is insufficient. Problem statements translate naturally from Lean to Dafny, as seen in Figure 1.
Problems span multiple mathematical domains including algebra, number theory, inequalities, combinatorics, and analysis. Each split (test and validation) contains 45 AMC problems, 15 AIME problems, and 20 IMO problems, with the remaining problems drawn from undergraduate mathematics courses. The benchmark tests both the ability to provide high-level proof strategies (lemma invocations, case splits, induction) and low-level proof details (algebraic manipulations, inequality chains).
Two supporting files provide mathematical infrastructure, though both remain work in progress. The definitions.dfy file (300 lines) axiomatizes core mathematical structures not native to Dafny: integers, rationals, reals, and complex numbers with their operations and properties. The library.dfy file (550 lines) contains 108 axiomatized lemmas encoding standard mathematical facts covering exponentials, logarithms, trigonometric functions, complex numbers, and number theory. While these could in principle be proven, we axiomatize them to focus the evaluation on proof synthesis with SMT automation. Rather than investing in deep library development, we set out on the goal of testing how well modern AI systems can develop proofs using Dafny’s Hoare-logic style specifications and SMT solver automation.
2.2. Definitions
function {:axiom} sum<T>(s: set<T>, f: T -> real): (p: real)
ensures s == {} ==> p == 0.0
ensures forall x | x in s :: p == f(x) + sum(s - {x}, f)
Figure 2. Example definition from definitions.dfy: sum replicating an inductive definition as postconditions.
function {:axiom} log(x: real): (l: real)
ensures x > 1.0 <==> l > 0.0
ensures 0.0 < x < 1.0 <==> l < 0.0
ensures x == 1.0 <==> l == 0.0
function {:axiom} logb(b: real, x: real): (l: real)
ensures x > 1.0 <==> l > 0.0
ensures 0.0 < x < 1.0 <==> l < 0.0
ensures x == 1.0 <==> l == 0.0
ensures b != 1.0 && b > 0.0 && x > 0.0 ==> logb(b, x) == log(x) / log(b)
Figure 3. Example definitions from definitions.dfy: log and logb specifying range behavior through postconditions.
The benchmark’s mathematical foundations are provided through definitions.dfy, which axiomatizes the minimal mathematical infrastructure needed to express all miniF2F problem statements in Dafny. Unlike Mathlib (mathlib), where mathematical objects are constructed from first principles, our axiomatization strategy reflects Dafny’s design philosophy of leveraging SMT solver capabilities for automated reasoning from a trusted library of definitions rather than requiring extensive proof libraries.
The definitions are organized across four numeric domains: integers, rationals, reals, and complex numbers. Each domain includes standard arithmetic operations and domain-specific functions. Integers support summation and product over finite sets, modular arithmetic, and divisibility predicates. Rationals are represented explicitly as numerator-denominator pairs with fraction arithmetic. Reals include transcendental functions (exponential, logarithm, trigonometric), power operations, and mathematical constants like $\pi$. Complex numbers provide field operations, norm functions, and complex exponential.
The axiomatization employs Dafny’s :axiom attribute to declare functions through their specifications rather than implementations. The specifications vary in complexity depending on the function’s nature. For transcendental functions like log (Figure 3), we axiomatize basic range properties—for instance, specifying that the logarithm is positive when the argument exceeds 1, negative when between 0 and 1, and zero at 1. For recursive functions like sum (Figure 2), we replicate inductive definitions as postconditions, specifying behavior on the empty set and on non-empty sets by removing one element. Number-theoretic functions (gcd, lcm, prime), combinatorial functions (choose, factorial), and utility functions (floor, ceil, abs) are similarly axiomatized. These contracts provide sufficient semantic information for Dafny’s SMT-backed automation to verify many problems with empty proofs. While the axiomatizations follow standard mathematical definitions and we are cautiously confident in their correctness, this pragmatic approach does come with the usual soundness considerations inherent to axiomatization.
2.3. Library
lemma {:axiom} logb_change_base(b1: real, b2: real, x: real)
requires b1 > 0.0 && b1 != 1.0
requires b2 > 0.0 && b2 != 1.0
requires x > 0.0
ensures logb(b1, x) == logb(b2, x) / logb(b2, b1)
Figure 4. Example lemma from library.dfy: the logarithm change of base formula.
Complementing the definitions, library.dfy provides 108 axiomatized lemmas encoding standard mathematical facts. The library was developed through a problem-driven approach: lemmas were identified by analyzing proof attempts to determine needed facts, then cross-checked against Mathlib (mathlib) theorems for soundness before integration. Rather than providing constructive proofs, we axiomatize these lemmas to enable problems to invoke established results without reproving them, mirroring mathematical practice where known theorems are cited rather than reconstructed. The lemmas span several mathematical domains:
  • Exponential/Logarithm (27 lemmas): Multiplication of exponentials, addition of logarithms, change of base formula (Figure 4)
  • Powers (8 lemmas): Natural and real exponentiation, power laws, square roots
  • Trigonometry (37 lemmas): Angle addition, periodicity, special values, Pythagorean identity
  • Complex Numbers (19 lemmas): Field axioms, norm properties, Euler’s formula
  • Number Theory (5 lemmas): GCD commutativity, GCD-LCM product formula
  • Analysis (4 lemmas): Limit uniqueness, continuity properties
  • Sequences (8 lemmas): Summation and product over sequences, set conversions
Each lemma is specified through preconditions and postconditions. For example, the change of base formula for logarithms (Figure 4) requires positive bases different from 1 and ensures the standard relationship between logarithms in different bases. Additionally, some mathematical properties—such as associativity and additivity—cannot be naturally encoded as postconditions on individual functions and must instead be stated as separate lemmas.
The library’s scope (550 lines) is intentionally minimal compared to Mathlib’s 2M lines, focusing on prerequisites for olympiad-level mathematics. This reflects a pragmatic balance: provide sufficient theory to express and solve miniF2F problems while testing how well AI systems can develop proofs from a compact foundation using SMT automation. The library is likely incomplete, and extending it with additional lemmas would likely improve evaluation results.
2.4. Validation
A critical component of our benchmark is validating that generated solutions adhere to the original problem statement without weakening it. This prevents solutions from strengthening preconditions, weakening postconditions, or introducing unsound axioms—issues we have observed in other benchmarks such as DafnyBench (loughridge2024dafnybenchbenchmarkformalsoftware), where weak evaluation scripts allowed such violations. Given a problem stated as a Dafny lemma, our validator rejects solutions that:
  • Contain warnings or errors during verification
  • Modify or remove original requires clauses
  • Weaken or remove original ensures clauses
  • Use the :axiom attribute to assume facts without proof
  • Use assume statements to bypass verification
Our validation pipeline operates in two stages. First, we invoke the Dafny verifier and parse its JSON output to detect verification failures. Second, we parse each problem lemma into its signature, preconditions, and postconditions using an extraction pipeline. The validator compares original and generated specifications: requires clauses must match exactly (no additions or removals), while ensures clauses must be a superset of the original. Here, adding postconditions is permitted because the strengthened lemma still implies the original statement—one can derive the original lemma from the stronger version. We additionally scan for :axiom attributes and assume statements. Critically, we distinguish verification diagnostics by source file—we allow warnings from library files (definitions.dfy, library.dfy) containing :axiom attributes, but any warnings or errors from the problem file itself trigger rejection. When validation fails, we return structured feedback indicating which clauses were modified or which unsound constructs were used, enabling models to self-correct in subsequent iterations.
3. Evaluation
We evaluate miniF2F-Dafny on baseline Dafny automation and 12 off-the-shelf LLMs to assess the effectiveness of auto-active verification for automated mathematical reasoning. Our evaluation first measures how many problems Dafny’s SMT-backed automation can solve without human intervention, then evaluates whether modern LLMs can provide proof hints to guide verification on the remaining problems.
3.1. Empty Proof Baseline
Dafny’s verifier, powered by Z3, verifies 99 of 244 test problems (40.6%) and 109 of 244 validation problems (44.7%) with empty proofs—requiring zero manual proof steps. We run each problem for 5 attempts with a 30-second timeout and use Dafny version 4.11.0. This baseline demonstrates the power of SMT-backed automation: problems that would require explicit proof terms in interactive theorem provers like Lean are automatically verified by Dafny. For example, IMO 1959 Problem 1 (Figure 1) requires a substantial proof in Lean but verifies with an empty proof in Dafny.
3.2. LLM-Guided Proofs
include "../definitions.dfy"
include "../library.dfy"
lemma imo_1964_p2(a: real, b: real, c: real)
requires 0.0 < a && 0.0 < b && 0.0 < c
requires c < a + b && b < a + c && a < b + c
ensures a*a*(b+c-a) + b*b*(c+a-b) + c*c*(a+b-c) <=
3.0*a*b*c
assert b + c - a > 0.0;
assert c + a - b > 0.0;
assert a + b - c > 0.0;
var lhs := a*a*(b+c-a) + b*b*(c+a-b) + c*c*(a+b-c);
calc {
lhs;
== a*a*b + a*a*c - a*a*a + b*b*c + b*b*a - b*b*b
+ c*c*a + c*c*b - c*c*c;
== a*a*b + a*a*c + b*b*c + b*b*a + c*c*a + c*c*b
- a*a*a - b*b*b - c*c*c;
var expr := a*a*a + b*b*b + c*c*c + 3.0*a*b*c
- a*a*b - a*a*c - b*b*a - b*b*c - c*c*a - c*c*b;
var sos1 := (a-b)*(a-b)*(a+b-c);
var sos2 := (b-c)*(b-c)*(b+c-a);
var sos3 := (c-a)*(c-a)*(c+a-b);
assert sos1 >= 0.0;
assert sos2 >= 0.0;
assert sos3 >= 0.0;
sos_identity(a, b, c);
assert expr >= 0.0;
assert 3.0*a*b*c - lhs >= 0.0;
lemma sos_identity(a: real, b: real, c: real)
requires 0.0 < a && 0.0 < b && 0.0 < c
ensures 2.0*(a*a*a + b*b*b + c*c*c + 3.0*a*b*c -
a*a*b - a*a*c - b*b*a - b*b*c - c*c*a - c*c*b) ==
(a-b)*(a-b)*(a+b-c) + (b-c)*(b-c)*(b+c-a) +
(c-a)*(c-a)*(c+a-b)
{ /* proof omitted for brevity */ }
Figure 5. IMO 1964 inequality problem solved by Claude Sonnet 4 using an auxiliary sum-of-squares identity lemma.
include "../definitions.dfy"
include "../library.dfy"
lemma amc12_2000_p6(p: nat, q: nat)
requires prime(p)
requires prime(q)
requires 4 <= p <= 18
requires 4 <= q <= 18
ensures p*q - (p+q) != 194
assert 195 % 2 == 1; // 195 is odd
assert 4 % 2 == 0; // 4 is even
assert 6 % 2 == 0; // 6 is even
assert 10 % 2 == 0; // 10 is even
assert 12 % 2 == 0; // 12 is even
assert 16 % 2 == 0; // 16 is even
Figure 6. AMC 12 prime product problem solved by Qwen3 Coder 30B using concise parity reasoning.
For problems where empty proofs fail, we evaluate 12 off-the-shelf LLMs on providing proof hints. Due to time and capacity constraints, some models are evaluated on a subset of problems; this evaluation is work in progress and will be updated with complete results.
Experimental Setup
Each problem receives up to $N=4$ generation attempts with $E=3$ error correction iterations after the initial attempt (temperature $T=0.5$, 8192 token limit per response). When verification fails, we extract error traces from Dafny’s output and append them to the conversation history for iterative refinement. We evaluate models via the AWS Bedrock API: Claude (Sonnet 3.7, Sonnet 4, Sonnet 4.5, Haiku 4.5), DeepSeek V3.1, Llama 4 Maverick 17B, GPT-OSS (20B, 120B), and Qwen 3 (32B, 235B MoE, Coder 30B, Coder 480B MoE). No models are fine-tuned.
Results
Table 1 presents pass@4 results on the test set. Claude Sonnet 4.5 achieves the highest performance at 55.7 %, followed by Claude Sonnet 3.7 (55.2 %) and Qwen 3 235B MoE (54.3 %). Several models cluster in the 43-50 % range. Models trained primarily on general code (DeepSeek V3.1, Llama 4 Maverick, GPT-OSS) show limited Dafny-specific knowledge, often confusing Dafny syntax with Lean. In contrast, larger models in the Claude and Qwen families demonstrate stronger familiarity with verification idioms in Dafny.
LLM-generated proofs exhibit a range of sophistication. At one end, Qwen3 Coder 30B solves an AMC 12 problem about prime products (Figure 6) with a concise parity argument: by asserting that 195 is odd while the relevant even numbers are even, enabling Dafny to automatically verify the result. At the other end, Claude Sonnet 4 solves a challenging IMO 1964 inequality problem (Figure 5) by introducing an auxiliary sum-of-squares identity lemma. The proof demonstrates sophisticated mathematical reasoning: it uses calc statements to algebraically manipulate the left-hand side, introduces three sum-of-squares terms that are non-negative by construction, and invokes the auxiliary lemma to establish the inequality. These examples illustrate that modern LLMs can generate both concise hints that leverage SMT automation and non-trivial proof strategies involving auxiliary lemmas and structured reasoning. Additional examples are provided in the appendix.
Error Analysis
Analysis of unverified problems reveals three main categories of difficulty:
  • Verification brittleness: Minor variations in assertion order or calc statement organization cause verification to fail.
  • Limited Dafny training data: Models struggle with language-specific idioms like calc statements and ghost variables, producing syntactically correct but semantically ineffective proofs.
  • Mathematical complexity: Problems require mathematical facts not present in the library, necessitating theory development from scratch.
Discussion
Our results demonstrate that Dafny’s SMT-backed automation provides a strong baseline for olympiad mathematics, with modern LLMs capable of providing effective proof hints to extend this further. Top models reach approximately 55% on the test set, representing over a 35% relative improvement over the 40% baseline. However, significant room for improvement remains, particularly in handling verification brittleness and expanding model familiarity with Dafny-specific proof patterns.
4. Related Work
4.1. Formal Mathematical Reasoning Benchmarks
Formal mathematical reasoning benchmarks provide automatic verification mechanisms through proof assistants, in contrast to informal benchmarks like MATH (hendrycks2021measuring) and GSM8K (cobbe2021training), which evaluate natural language mathematical reasoning without formal guarantees of correctness.
The miniF2F benchmark (minif2f) is a Lean benchmark consisting of 488 mathematical problems drawn from the AIME, AMC, and IMO, as well as material from high school and undergraduate mathematics courses, split into test and validation sets of 244 problems each. It has been translated to Isabelle, HOL Light, and Metamath, all of which are interactive theorem provers requiring explicit proof terms.
Other competition-style benchmarks include FIMO (liu2023fimo), which contains Lean formalizations of IMO shortlist problems and PutnamBench (tsoukalas2024putnambench), which contains Lean problems from the William Lowell Putnam Mathematical Competition and represents a significantly higher level of difficulty. ProofNet (azerbayev2023proofnet) focuses on Lean exercises from undergraduate mathematics textbooks. LeanDojo (yang2023leandojo) provides a dataset derived from Lean’s mathlib library.
MiniF2F remains the most widely adopted benchmark due to its manageable size, diverse problem coverage, and multi-language implementations.
4.2. AI for Interactive Theorem Proving
Whole-proof approaches generate complete proofs that are iteratively refined until verified. GPT-f (polu2020generative) pioneered this in Metamath, followed by FMS-CL (polu2022formalmathematicsstatementcurriculum) in Lean. More recent whole-proof systems include DeepSeek-Prover (xin2024deepseek), Seed-Prover (seed-prover), and Kimina-Prover (wang2025kiminaproverpreviewlargeformal).
Step-wise approaches construct proofs incrementally using tree search, including HTPS (lample2022hypertree) (Lean and Metamath) and LLEMMA (azerbayev2023llemma) (Lean). Leadership on benchmarks has alternated between whole-proof and step-wise paradigms.
Hybrid systems combine informal reasoning with formal verification. DSP (jiang2022draft) and LEGO-Prover (wang2023lego) work in Isabelle, using natural language proofs converted to formal sketches.
Agentic frameworks orchestrate multiple components for proof generation. Systems include COPRA (thakur2023context), Prover-Agent (baba2025prover), ProofCompass (wischermann2025proofcompass), and HILBERT (hilbert) in Lean.
Domain-specific approaches like AlphaGeometry (trinh2024solving) target specific problem types, such as IMO geometry problems.
Success rates have improved dramatically recently, with HILBERT achieving 99.2% on miniF2F and 70.0% on PutnamBench. Multiple systems achieved gold medal performance on IMO 2025, including Seed-Prover and Aristotle (achim2025aristotle) with formal solutions, and systems from Google DeepMind and OpenAI with natural language solutions.
5. Future Work
Dafny-specific training
Current models exhibit limited exposure to Dafny syntax and verification idioms, often conflating Dafny with interactive theorem provers. A key limitation is their inability to judge Z3’s capacity to automatically discharge subgoals. Pre-training on curated Dafny corpora and fine-tuning on verification tasks could improve familiarity with auto-active verification patterns and teach effective use of calc statements, assertion placement, and ghost variables.
Agentic architectures
Analogous to Lean-based systems like Aristotle (achim2025aristotle), Seed-Prover (seed-prover), and HILBERT (hilbert), Dafny-specific agentic frameworks could orchestrate proof search, lemma synthesis, and iterative refinement, exploiting the synergy between program synthesis and formal verification.
Learned lemma libraries
LEGO-Prover (wang2023lego) shows how models can extract, generalize, and cache successful proof strategies as reusable lemmas. Adapting this to Dafny could enable cumulative learning across problems, mirroring human mathematical practice.
6. Conclusion
miniF2F-Dafny represents the first exploration of auto-active verification for pure mathematical reasoning, a domain traditionally dominated by interactive theorem provers. Our results demonstrate an effective division of labor: LLMs provide high-level guidance while SMT-backed automation handles low-level details. This paradigm leverages the complementary strengths of automated reasoning and modern language models. Looking forward, we anticipate convergence between auto-active verification and interactive theorem proving—recent developments like Lean’s grind tactic already demonstrate this trend. Our work suggests a promising path toward accessible, AI-assisted formal verification that exploits this synergy.
Appendix A Example Solutions
In this section, we present some representative proof solutions generated by LLMs to illustrate the range of proof strategies employed in miniF2F-Dafny. We choose examples demonstrating sophisticated mathematical reasoning with auxiliary lemmas to show that the combination of Dafny’s verification combined with precise LLM output can produce complete solutions to IMO-level mathematical problems.
A.1. Complete Proof of imo_1964_p2
The solution to the IMO 1964 problem 2 (shown truncated in Section 4) synthesized by Claude Sonnet 4 demonstrates a sophisticated proof strategy using sum-of-squares decomposition. The proof establishes a triangle inequality by showing that $3abc-\text{LHS}$ can be expressed as a sum of squared terms, each multiplied by positive factors derived from the triangle inequalities. This requires proving an auxiliary algebraic identity through systematic expansion and term collection. Using the auxiliary facts, a complete proof is verified by Dafny’s SMT solver based backend.
include "../definitions.dfy"
include "../library.dfy"
lemma imo_1964_p2(a: real, b: real, c: real)
requires 0.0 < a
requires 0.0 < b
requires 0.0 < c
requires c < a + b
requires b < a + c
requires a < b + c
ensures a*a*(b+c-a) + b*b*(c+a-b) + c*c*(a+b-c) <= 3.0*a*b*c
// First, note that the triangle inequalities ensure all terms are positive
assert b + c - a > 0.0;
assert c + a - b > 0.0;
assert a + b - c > 0.0;
// We’$l$ prove that 3abc - LHS >= 0 by showing it equals a sum of squares
var lhs := a*a*(b+c-a) + b*b*(c+a-b) + c*c*(a+b-c);
// Expand the LHS
calc {
lhs;
== a*a*(b+c-a) + b*b*(c+a-b) + c*c*(a+b-c);
== a*a*b + a*a*c - a*a*a + b*b*c + b*b*a - b*b*b + c*c*a + c*c*b - c*c*c;
== a*a*b + a*a*c + b*b*c + b*b*a + c*c*a + c*c*b - a*a*a - b*b*b - c*c*c;
// Now compute 3abc - lhs
calc {
3.0*a*b*c - lhs;
== 3.0*a*b*c - (a*a*b + a*a*c + b*b*c + b*b*a + c*c*a + c*c*b - a*a*a - b*b*b - c*c*c);
== 3.0*a*b*c - a*a*b - a*a*c - b*b*c - b*b*a - c*c*a - c*c*b + a*a*a + b*b*b + c*c*c;
== a*a*a + b*b*b + c*c*c + 3.0*a*b*c - a*a*b - a*a*c - b*b*a - b*b*c - c*c*a - c*c*b;
// We’$l$ show this equals (1/2) times a sum of squares
var expr := a*a*a + b*b*b + c*c*c + 3.0*a*b*c - a*a*b - a*a*c - b*b*a - b*b*c - c*c*a - c*c*b;
// Define the sum of squares terms with explicit non-negativity proofs
var diff_ab := a - b;
var diff_bc := b - c;
var diff_ca := c - a;
// Squares are always non-negative
assert diff_ab * diff_ab >= 0.0;
assert diff_bc * diff_bc >= 0.0;
assert diff_ca * diff_ca >= 0.0;
// Triangle inequality factors are positive
assert a + b - c > 0.0;
assert b + c - a > 0.0;
assert c + a - b > 0.0;
var sos1 := diff_ab * diff_ab * (a + b - c);
var sos2 := diff_bc * diff_bc * (b + c - a);
var sos3 := diff_ca * diff_ca * (c + a - b);
// Each SOS term is non-negative (product of non-negative terms)
assert sos1 >= 0.0;
assert sos2 >= 0.0;
assert sos3 >= 0.0;
// Prove the key identity: 2*expr = sos1 + sos2 + sos3
sos_identity(a, b, c);
// Therefore expr = (1/2)(sos1 + sos2 + sos3) >= 0
assert expr >= 0.0;
assert 3.0*a*b*c - lhs >= 0.0;
lemma sos_identity(a: real, b: real, c: real)
requires 0.0 < a
requires 0.0 < b
requires 0.0 < c
ensures 2.0*(a*a*a + b*b*b + c*c*c + 3.0*a*b*c - a*a*b - a*a*c - b*b*a - b*b*c - c*c*a - c*c*b)
== (a-b)*(a-b)*(a+b-c) + (b-c)*(b-c)*(b+c-a) + (c-a)*(c-a)*(c+a-b)
var lhs := 2.0*(a*a*a + b*b*b + c*c*c + 3.0*a*b*c - a*a*b - a*a*c - b*b*a - b*b*c - c*c*a - c*c*b);
var rhs := (a-b)*(a-b)*(a+b-c) + (b-c)*(b-c)*(b+c-a) + (c-a)*(c-a)*(c+a-b);
// Expand each term on the right-hand side systematically
var term1_expanded := expand_first_term(a, b, c);
var term2_expanded := expand_second_term(a, b, c);
var term3_expanded := expand_third_term(a, b, c);
// Sum all three expanded terms
calc {
rhs;
== term1_expanded + term2_expanded + term3_expanded;
== (a*a*a - a*a*b - a*b*b - a*a*c + 2.0*a*b*c + b*b*b - b*b*c) +
(b*b*b - b*b*c - a*b*b - b*c*c + 2.0*a*b*c + c*c*c - a*c*c) +
(c*c*c - a*c*c - b*c*c - a*a*c + 2.0*a*b*c + a*a*a - a*a*b);
// Collect like terms carefully
calc {
rhs;
== 2.0*a*a*a + 2.0*b*b*b + 2.0*c*c*c + 6.0*a*b*c - 2.0*a*a*b - 2.0*a*b*b - 2.0*a*a*c - 2.0*a*c*c - 2.0*b*b*c - 2.0*b*c*c;
// Use commutativity of multiplication
assert a*b*b == b*b*a;
assert a*c*c == c*c*a;
assert b*c*c == c*c*b;
calc {
rhs;
== 2.0*a*a*a + 2.0*b*b*b + 2.0*c*c*c + 6.0*a*b*c - 2.0*a*a*b - 2.0*b*b*a - 2.0*a*a*c - 2.0*c*c*a - 2.0*b*b*c - 2.0*c*c*b;
== 2.0*(a*a*a + b*b*b + c*c*c + 3.0*a*b*c - a*a*b - b*b*a - a*a*c - c*c*a - b*b*c - c*c*b);
== lhs;
function expand_first_term(a: real, b: real, c: real): real
a*a*a - a*a*b - a*b*b - a*a*c + 2.0*a*b*c + b*b*b - b*b*c
lemma expand_first_term_correct(a: real, b: real, c: real)
ensures (a-b)*(a-b)*(a+b-c) == expand_first_term(a, b, c)
calc {
(a-b)*(a-b)*(a+b-c);
== (a*a - 2.0*a*b + b*b)*(a+b-c);
== a*a*a + a*a*b - a*a*c - 2.0*a*a*b - 2.0*a*b*b + 2.0*a*b*c + b*b*a + b*b*b - b*b*c;
== a*a*a - a*a*b - a*b*b - a*a*c + 2.0*a*b*c + b*b*b - b*b*c;
== expand_first_term(a, b, c);
function expand_second_term(a: real, b: real, c: real): real
b*b*b - b*b*c - a*b*b - b*c*c + 2.0*a*b*c + c*c*c - a*c*c
lemma expand_second_term_correct(a: real, b: real, c: real)
ensures (b-c)*(b-c)*(b+c-a) == expand_second_term(a, b, c)
calc {
(b-c)*(b-c)*(b+c-a);
== (b*b - 2.0*b*c + c*c)*(b+c-a);
== b*b*b + b*b*c - a*b*b - 2.0*b*b*c - 2.0*b*c*c + 2.0*a*b*c + b*c*c + c*c*c - a*c*c;
== b*b*b - b*b*c - a*b*b - b*c*c + 2.0*a*b*c + c*c*c - a*c*c;
== expand_second_term(a, b, c);
function expand_third_term(a: real, b: real, c: real): real
c*c*c - a*c*c - b*c*c - a*a*c + 2.0*a*b*c + a*a*a - a*a*b
lemma expand_third_term_correct(a: real, b: real, c: real)
ensures (c-a)*(c-a)*(c+a-b) == expand_third_term(a, b, c)
calc {
(c-a)*(c-a)*(c+a-b);
== (c*c - 2.0*c*a + a*a)*(c+a-b);
== c*c*c + a*c*c - b*c*c - 2.0*a*c*c - 2.0*a*a*c + 2.0*a*b*c + a*a*c + a*a*a - a*a*b;
== c*c*c - a*c*c - b*c*c - a*a*c + 2.0*a*b*c + a*a*a - a*a*b;
== expand_third_term(a, b, c);
Listing 1: Complete proof of imo_1964_p2 using a sum-of-squares identity.
A.2. Complete Proof of algebra_amgm_sumasqdivbgeqsuma
The algebra_amgm_sumasqdivbgeqsuma problem requires a proof of the algebraic inequality: $\frac{a^{2}}{b}+\frac{b^{2}}{c}+\frac{c^{2}}{d}+\frac{d^{2}}{a}\geq a+b+c+d$ for positive reals $a,b,c,d$. The solution demonstrates sophisticated proof engineering by reducing the problem to the Cauchy-Schwarz inequality. The proof, synthesized by Claude Sonnet 4, introduces auxiliary lemmas to establish the Cauchy-Schwarz inequality for four terms and uses a variable substitution ($x_{i}=\frac{a}{\sqrt{b}}$, $y_{i}=\sqrt{b}$, …) to transform the target inequality into a form where Cauchy-Schwarz applies directly. This exemplifies how modern LLMs can identify complex proof strategies in Dafny and structure complex arguments through helper lemmas.
include "../definitions.dfy"
include "../library.dfy"
lemma cauchy_schwarz_four_terms(x1: real, x2: real, x3: real, x4: real, y1: real, y2: real, y3: real, y4: real)
ensures (x1*y1 + x2*y2 + x3*y3 + x4*y4) * (x1*y1 + x2*y2 + x3*y3 + x4*y4) <=
(x1*x1 + x2*x2 + x3*x3 + x4*x4) * (y1*y1 + y2*y2 + y3*y3 + y4*y4)
var lhs := (x1*y1 + x2*y2 + x3*y3 + x4*y4) * (x1*y1 + x2*y2 + x3*y3 + x4*y4);
var rhs := (x1*x1 + x2*x2 + x3*x3 + x4*x4) * (y1*y1 + y2*y2 + y3*y3 + y4*y4);
var diff := rhs - lhs;
var sum_of_squares := (x1*y2 - x2*y1)*(x1*y2 - x2*y1) +
(x1*y3 - x3*y1)*(x1*y3 - x3*y1) +
(x1*y4 - x4*y1)*(x1*y4 - x4*y1) +
(x2*y3 - x3*y2)*(x2*y3 - x3*y2) +
(x2*y4 - x4*y2)*(x2*y4 - x4*y2) +
(x3*y4 - x4*y3)*(x3*y4 - x4*y3);
assert sum_of_squares >= 0.0;
lemma division_by_sqrt_product(x: real, y: real)
requires x > 0.0
requires y > 0.0
ensures x / (sqrt(y) * sqrt(y)) == x / y
assert sqrt(y) * sqrt(y) == y;
lemma algebra_amgm_sumasqdivbgeqsuma(a: real, b: real, c: real, d: real)
requires 0.0 < a
requires 0.0 < b
requires 0.0 < c
requires 0.0 < d
ensures (a*a)/b + (b*b)/c + (c*c)/d + (d*d)/a >= a + b + c + d
var x1 := a / sqrt(b);
var x2 := b / sqrt(c);
var x3 := c / sqrt(d);
var x4 := d / sqrt(a);
var y1 := sqrt(b);
var y2 := sqrt(c);
var y3 := sqrt(d);
var y4 := sqrt(a);
assert sqrt(a) > 0.0;
assert sqrt(b) > 0.0;
assert sqrt(c) > 0.0;
assert sqrt(d) > 0.0;
division_by_sqrt_product(a*a, b);
division_by_sqrt_product(b*b, c);
division_by_sqrt_product(c*c, d);
division_by_sqrt_product(d*d, a);
cauchy_schwarz_four_terms(x1, x2, x3, x4, y1, y2, y3, y4);
assert x1*x1 == (a/sqrt(b)) * (a/sqrt(b));
assert x1*x1 == (a*a) / (sqrt(b) * sqrt(b));
assert x1*x1 == (a*a) / b;
assert x2*x2 == (b*b) / c;
assert x3*x3 == (c*c) / d;
assert x4*x4 == (d*d) / a;
var sum_x_sq := x1*x1 + x2*x2 + x3*x3 + x4*x4;
var target_lhs := (a*a)/b + (b*b)/c + (c*c)/d + (d*d)/a;
assert sum_x_sq == target_lhs;
assert x1*y1 == (a/sqrt(b)) * sqrt(b) == a;
assert x2*y2 == (b/sqrt(c)) * sqrt(c) == b;
assert x3*y3 == (c/sqrt(d)) * sqrt(d) == c;
assert x4*y4 == (d/sqrt(a)) * sqrt(a) == d;
var sum_xy := x1*y1 + x2*y2 + x3*y3 + x4*y4;
assert sum_xy == a + b + c + d;
assert y1*y1 == sqrt(b) * sqrt(b) == b;
assert y2*y2 == sqrt(c) * sqrt(c) == c;
assert y3*y3 == sqrt(d) * sqrt(d) == d;
assert y4*y4 == sqrt(a) * sqrt(a) == a;
var sum_y_sq := y1*y1 + y2*y2 + y3*y3 + y4*y4;
assert sum_y_sq == a + b + c + d;
assert sum_xy * sum_xy <= sum_x_sq * sum_y_sq;
assert (a + b + c + d) * (a + b + c + d) <= target_lhs * (a + b + c + d);
assert a + b + c + d > 0.0;
assert (a + b + c + d) <= target_lhs;
Listing 2:
Complete proof of algebra_amgm_sumasqdivbgeqsuma using the Cauchy-Schwarz inequality.
Appendix B Prompts
B.1. Initial Proof Synthesis Prompt
We craft an initial model prompt to establish the task context and constraints that we present below. It emphasizes preserving problem specifications, prohibits unsound constructs (axioms, assumptions), and encourages strategic use of Dafny idioms like calc statements and assertions. The prompt explicitly instructs models to prove any mathematical results they invoke, preventing reliance on unstated ”classical theorems.” This design reflects our goal of testing proof synthesis rather than library lookup.
      # ROLE #    You are an expert Dafny programmer specializing in formal    verification, proof construction, mathematical reasoning and    proof hint infilling.        # OBJECTIVE #    Complete the proof body of the given incomplete Dafny lemma    to make it verify successfully. The lemma currently has an    empty body ‘{}‘ that you must replace with a complete,    verifying proof.        # AVAILABLE RESOURCES #        ## definitions.dfy file content:    {definitions_file_content}        ## library.dfy file content:    {library_file_content}        # TASK REQUIREMENTS #        You will receive an incomplete Dafny file containing a lemma    with:    - A complete signature (name and parameters)    - Complete ‘requires‘ clauses (preconditions)    - Complete ‘ensures‘ clauses (postconditions)    - An empty body ‘{}‘ that needs to be filled        Your task is to replace the empty body with a complete proof    that satisfies all postconditions.        # CRITICAL RULES #        1. PRESERVE SPECIFICATION: You MUST preserve the exact lemma     signature and all ‘requires‘ clauses. Do NOT add, remove,     or modify any ‘requires‘ clauses.        2. ENSURES CLAUSES: The ‘ensures‘ clauses MUST all be     preserved. You MAY add additional ‘ensures‘ clauses if     they may aid with verification but this is not encouraged.        3. COMPLETE IMPLEMENTATION: Provide a complete proof body.     This may include:     - Assertions to guide the verifier     - Calc statements for step-by-step reasoning     - Calls to helper lemmas (from library.dfy or your own)     - Loop invariants     - Case analysis or proof by contradiction        4. USE AVAILABLE RESOURCES: Leverage functions and lemmas     from definitions.dfy and library.dfy without redefining     them. They are already imported via     ‘include "../definitions.dfy"‘ and     ‘include "../library.dfy"‘.        5. DEFINE HELPERS IF NEEDED: If you need helper lemmas or     functions not available in the provided files, define     them within your solution.        6. NO ASSUME OR AXIOM ATTRIBUTES: You may NOT use the     ‘assume‘ statement or ‘{:axiom}‘ attribute anywhere in     your solution. All auxiliary lemmas MUST include complete     proofs.        7. PROVE ALL CLAIMS: If you reference mathematical results,     theorems, or inequalities (e.g. Cauchy-Schwarz inequality,     AM-GM, Schur’s Inequality etc.) that are not present in     the definitions.dfy or library.dfy files, you MUST prove     them yourself within your solution. Do NOT rely on     "classical results" or "known theorems" without providing     a complete formal proof in Dafny.        8. EXPLICIT TRIGGERS FOR QUANTIFIERS: You MUST add explicit     triggers to all quantifiers (forall/exists) where they     are not automatically synthesized by Dafny. Use the     ‘{:trigger}‘ attribute to specify triggers that help     guide the verifier’s instantiation of quantified variables.        # PROOF STRATEGIES #        Consider these approaches when constructing your proof:    - Direct proof: Use assertions and calc statements to show     the postcondition follows from preconditions    - Calc statements: Use ‘calc‘ blocks for step-by-step     algebraic calculations and equational reasoning - this is     especially useful for mathematical proofs involving     arithmetic, logarithms, or algebraic manipulations    - Case analysis: Split the proof into cases based on     conditions    - Induction: For properties involving natural numbers or     recursive structures    - Contradiction: Assume the negation and derive a     contradiction    - Helper lemmas: Break complex proofs into smaller,     manageable pieces    - Verifier hints: Add strategic assertions to guide Dafny’s     automated reasoning        The complete file must include:    - The original ‘include‘ statement    - The complete lemma with your proof body    - Any helper lemmas or functions you define        # INPUT FORMAT    ‘‘‘dafny    <INCOMPLETE_FILE_CONTENT>    ‘‘‘        # RESPONSE FORMAT #        Reason step-by-step first and then, provide the complete    Dafny file in this exact format:    ‘‘‘dafny    <COMPLETE_FILE_CONTENT>    ‘‘‘        ---        ‘‘‘dafny    {incomplete_dafny_file_content}    ‘‘‘       Listing 3. System prompt for initial proof generation.
B.2. Error Correction Proof Synthesis Prompt
When verification fails, the error correction prompt presented below provides structured feedback including Dafny’s error messages and a reminder of the original specification. This enables iterative refinement while maintaining specification adherence. We keep the error-correction prompt brief for the purposes of reducing tokens used and maximizing information efficiency by avoiding repetition.
      Your code failed Dafny verification with the following errors:        {errors}        Reminder - the original lemma you need to complete is:    ‘‘‘dafny    {incomplete_dafny_file_content}    ‘‘‘        Please fix these errors and provide the complete corrected    Dafny file. CRITICAL:    - You MUST preserve the exact lemma signature and all     ‘requires‘ clauses from the original lemma above    - You MUST keep all ‘ensures‘ clauses from the original     lemma above    - Only the proof body should be modified - do NOT change     the specification    - Do NOT use ‘assume‘ statements or ‘{:axiom}‘ attributes    - Add explicit triggers to quantifiers where needed        Provide the corrected code in this format:    ‘‘‘dafny    <COMPLETE_CORRECTED_FILE_CONTENT>    ‘‘‘       Listing 4. System prompt for initial proof generation.